OpenStudy (anonymous):

find the sum : Σ(r=1 to n-1) {[r+1]^3-1}

OpenStudy (anonymous):

*[(r+1)^3 - 1 ]

OpenStudy (aaronandyson):

1+1^3-1

OpenStudy (aaronandyson):

2^3-1

OpenStudy (aaronandyson):

3-1 = 2

OpenStudy (aaronandyson):

2^2 = ?

OpenStudy (aaronandyson):

2^2 = 2 times 2 = ?

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

answer is in form of n

OpenStudy (michele_laino):

Hint: we have: \[{\left( {r + 1} \right)^3} - 1 = {r^3} + 3{r^2} + 3r\] so we can split your sum as follows: \[\sum\limits_1^{n - 1} {\left[ {{{\left( {r + 1} \right)}^3} - 1} \right]} = \sum\limits_1^{n - 1} {{r^3}} + 3\sum\limits_1^{n - 1} {{r^2}} + 3\sum\limits_1^{n - 1} r \]

OpenStudy (anonymous):

this one one will time taking there is one way to simplify that into r^3 -1 only